(41)Suppose the market value of prostitution in Sweden, a country where prostitution is a legalized is $75 billion a year, a macroeconomist who says that the GDP of her country will increase by $75, 000,000,000 a year is:

(a)Correct

(b)Incorrect

(c)(a) or (b) above

(d)More information needed to respond

                                                               Table#1-National Income Accounts

National Income Accounts

                                             $Billions

                                                  

Consumption expenditures

$1855

Interest

190

Corporate profit

150

Depreciation

200

Rent

70

Gross domestic investment

1250

Compensation of employees

830

Federal government expenditures

530

State government expenditures

235

Local government expenditures

140

Exports

220

Imports

270

(42)The GDP in Table#1 is equal to:           

(a)$3960b

(b)$5430b

(c)$3420b

(d)$6165b

(43)In Table#1, NDP equals:

(a)$2756b

(b)$3760b

(c)$4320b

(d)$5875b

(44)Which of the following represents the percent share of state government expenditures as a share of aggregate government expenditures in Table#1 above?

(a)7 percent

(b)16 percent

(c)26 percent

(d)None of the above

(45)Which of the following is the monetary value of capital consumption allowances for the economy represented in Table#1 above?

(a)$150b

(b)$1855b

(c)$830b

(d)None of the above
Answer #1

(41) (a)

Illegal goods/services are excluded from GDP. So when an illegal service is legalized, its value will increase by amount of its reported market value.

(42) (a)

GDP = consumption + gross domestic investment + Federal government expenditures + State government expenditures + Loal government expenditures + exports - imports

= 1855 + 1250 + 530 + 235 + 140 + 220 - 270

= 3960

(43) (b)

NDP = GDP - depreciation = 3960 - 200 = 3760

(44) (c)

% share = 235 / (530 + 235 + 140) = 235 / 905 = 0.26 = 26%

(45) (d)

Capital consumption allowance = depreciation = 200